Roni is 30 years old. Grace is 22 years old. How many years ago was Roni twice as old as Grace?
Found 3 solutions by josgarithmetic, Alan3354, ikleyn:
Answer by josgarithmetic(39618)
Answer by Alan3354(69443)

The difference is 8 years.
---> 8 & 16

Answer by ikleyn(52790)

.

Let x be "x years ago".


Then you have this equation


    30-x = 2(22-x).


Simplify and solve for x.


    30 - x = 44 - 2x

    2x - x = 44 - 30

     x     = 14


ANSWER.  14 years ago.


CHECK.  14 years ago Toni was 30-14 = 16 years old; Grace was  22-14 = 8.

        16 = 2*8.    ! Correct !

Ignore any other answer.